Graphing quadratic functions:
y= x^2+16x+29
I need the vertex and the direction
Thank you so much!

Answer by richwmiller(17219)

-b/2a=-16/2=-8 x of the vertex =-8
Plug in -8 for x and find y=-35
(-8,-35) vertex
It opens up
The vertex is the minimum.